OpenStudy (anonymous):

PLEASE HELP!!! Solve 7x - 2y = -3 14x + y = 14

OpenStudy (anonymous):

a) (4, 5 over 7) b) (4, 7 over 5) c) (5 over 7, 4) d) (7 over 5, 4)

OpenStudy (anonymous):

Lemme get dem answer choices !

jhonyy9 (jhonyy9):

multiplie second by 2 and will get 7x-2y=-3 28x +2y=28 -------------- 35x 0 = 25 x= 25/35 = 5/7 x=5/7 7*5/7 -2y =-3 5-2y =-3 -2y = -8 y= 4 x= 5/7 y= 4

jhonyy9 (jhonyy9):

a. look right